# Efficient Uncertainty Propagation in Bayesian Two-Step Procedures

### v1

@@ 2.3 Iterative Method
Input: Variable of interest $\theta\sim p(\theta\mid C)$, underlying variable $\tau\sim p(\tau\mid C)$
1 Draw $m$ samples $\tau^{(1)},\dots,\tau^{(m)}\sim p(\tau\mid C)$.
2 Set index set $I=\{1,\dots,m\}$.
3 while $|I|>0$ do
4       Select representative value $\tau^{*}\in\{\tau^{(i)}\}_{i\in I}$ and run MCMC to access the posterior distribution through samples $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p(\theta\mid\tau^{*})$.
5       Update index set $I$ such that $\tau^{*}\notin\{\tau^{(i)}\}_{i\in I}$
6       for $i\in I$ do
7             Run PSIS with target $p(\theta\mid\tau^{(i)})$ and proposal $p(\theta\mid\tau^{*})$ and gain importance weights and metric $\hat{k}_{i}$.
8             if $\hat{k}_{i}<0.7$ then
9                   Use calculated importance weights and draws $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p(\theta\mid\tau^{*})$ for importance resampling and save the resampled draws as posterior draws corresponding to $p(\theta\mid\tau^{(i)})$.
10                   Remove $i$ from $I$: $I\leftarrow I\setminus i$.
11
12            else
13                   Run IWMM with with target $p(\theta\mid\tau^{(i)})$ and proposal $p(\theta\mid\tau^{*})$ and gain importance weights and metric $\hat{k}_{\text{MM},i}$.
14                   if $\hat{k}_{\text{MM},i}<0.7$ then
15                         Use calculated importance weights and transformed draws $\{\breve{\theta}^{(s)}_{*}\}_{s=1}^{S}$ for importance resampling and save the resampled draws as posterior draws corresponding to $p(\theta\mid\tau^{(i)})$.
16                         Remove $i$ from $I$: $I\leftarrow I\setminus i$.
17
18
19
20
21Calculate the marginalized posterior via Monte Carlo Estimate $p(\theta\mid C)\approx\frac{1}{m}\sum_{i=1}^{m}p\left(\theta\mid\tau^{(i)}\right)$.
@@ 5       Update index set $I$ such that $\tau^{*}\notin\{\tau^{(i)}\}_{i\in I}$
Algorithm 1 provides accurate posterior draws through MCMC or resampled posterior draws through PSIS or IWMM for all $\{\tau^{(i)}\}_{i=1}^{m}$, corresponding to the desired distribution $p(\theta\mid\tau^{(i)})$. This allows the calculation of the Monte Carlo estimate from Equation (2), enabling us to approximate the marginalized posterior distribution $p(\theta\mid C)$. Since each posterior approximation used in the calculation is either based on computationally expensive MCMC or on PSIS/IWMM, where the quality of the approximation is ensured by accepting the resampling only when $\hat{k}$ or $\hat{k}_{\rm MM}$ is below the threshold, the accuracy of the estimate is maintained. The algorithm improves efficiency compared to running MCMC for all $m$ values separately, as each iteration (cycles of the while-loop in row 1) requires only one MCMC run for the proposal distribution, reducing the number of MCMC runs significantly. The improvement in efficiency is tested through extensive simulation studies in Section 3.
@@ 2.4 Special Case: Missing Data Problem
We now consider the setting in which the observed dataset $D_{\rm obs}=(y_{\rm obs},X_{\rm obs})$ contains missing values, potentially affecting both the target variable and the predictors. While there are straightforward approaches to handling missing data – such as listwise or pairwise deletion, and mean imputation [20] – these methods fail to propagate the uncertainty introduced by the missing values into the parameter estimates. A more principled approach that allows for uncertainty propagation is Multivariate Imputation using Chained Equations (MICE) [21]. MICE is an iterative algorithm that imputes the missing values in the dataset by generating $m$ completed versions of the data, each containing potentially different imputed values at the locations of the missing entries [23]. These multiple imputations help capture the uncertainty associated with the missing values in the original dataset $D_{\rm obs}$. We define the $m$ imputed datasets as $D^{(1)},\dots,D^{(m)}$, with observations in each dataset $D^{(i)}$ indexed by the superscript $(i)$ and denote the imputation process performed by MICE via the distribution $p_{\text{MICE}}(D\mid D_{\rm obs})$. Our goal is to characterize the marginalized posterior distribution of the model parameters $\theta$, which is obtained by integrating over the uncertainty in the imputed datasets:
[…]
For this problem setting, our method presented in Section 2.3 can be applied, with the context $C$ being the observed dataset $D_{\rm obs}$. A structural visualization of the iterative method (cf. Algorithm 1) is illustrated in Figure 3 for the special case of working with missing data.
@@
Calculation of Importance Ratios The specific case of handling missing data offers additional advantages that allow our algorithm to operate even more efficiently. In particular, the computation of importance ratios for PSIS and IWMM can be further optimized in this setting – an improvement not necessarily feasible in the more general iterative approach described in Section 2.3. This efficiency gain stems from the structural similarity among the imputed datasets $\{D^{(i)}\}_{i\in I}$, which differ only in the entries corresponding to originally missing values. We denote by $J^{*}\subset\{1,\dots,N\}$ the index set of observations where these differences occur.
@@
$r_{i}(\theta_{*}^{(s)})=\frac{p(\theta_{*}^{(s)}\mid D^{(i)})}{p(\theta_{*}^{(s)}\mid D^{*})}=\frac{p(\theta_{*}^{(s)}\mid y^{(i)},X^{(i)})}{p(\theta_{*}^{(s)}\mid y^{*},X^{*})}\propto\frac{p(y^{(i)}\mid\theta_{*}^{(s)},X^{(i)})p(\theta_{*}^{(s)})}{p(y^{*}\mid\theta_{*}^{(s)},X^{*})p(\theta_{*}^{(s)})}=\frac{p(y^{(i)}\mid\theta_{*}^{(s)},X^{(i)})}{p(y^{*}\mid\theta_{*}^{(s)},X^{*})}\,,s=1,\dots,S.$ (20)
@@
$r_{i}(\theta_{*}^{(s)})=\frac{\prod_{n=1}^{N}p(y_{n}^{(i)}\mid\theta_{*}^{(s)},X^{(i)})}{\prod_{n=1}^{N}p(y_{n}^{*}\mid\theta_{*}^{(s)},X^{*})}=\frac{\prod_{n\in J^{*}}p(y_{n}^{(i)}\mid\theta_{*}^{(s)},X^{(i)})}{\prod_{n\in J^{*}}p(y_{n}^{*}\mid\theta_{*}^{(s)},X^{*})}\,,s=1,\dots,S,$ (21)
because $p(y_{n}^{(i)}\mid\theta_{*}^{(s)},X^{(i)})=p(y_{n}^{*}\mid\theta_{*}^{(s)},X^{*})$ for all $n\notin J^{*}$. This reduction allows us to evaluate only the likelihood terms that correspond to differing rows, decreasing the number of pointwise likelihood evaluations by a factor of $|J^{*}|/N$, where $|J^{*}|$ corresponds to the number of rows with missing values. This optimization significantly reduces the computational effort required to compute the importance ratios for PSIS and IWMM, thereby making the overall workflow more efficient.
@@ 2.5 Special Case: Surrogate Modeling
Inference Step Next, we consider a new set of observations – typically real-world measurements – denoted by ${D_{I}=\{y_{I}^{(n)}\}_{n=1}^{N_{I}}}$. We assume that the (implicit) data-generating process underlying these observations is well captured by the simulator $\mathcal{M}$. In this context, the measurement data $D_{I}$ is thought to be generated by the simulator using some unknown input parameter $\theta_{I}$:
@@
[…]
@@
The selected draws should be representative in the sense that their corresponding posterior distributions collectively cover the diversity of possible posterior shapes – or, in the case of selecting a single draw, that the draw is central in terms of its posterior characteristics.
A straightforward approach is to apply clustering algorithms such as $k$-means [48] or $k$-medoids [38] in the space of the draws $\tau$. However, we cannot be sure that the similarity between surrogate parameter draws $\tau^{(i)}$ and $\tau^{(j)}$ does reliably reflect the similarity of their respective posterior distributions, $p(\theta\mid\tau^{(i)},D_{I})$ and $p(\theta\mid\tau^{(j)},D_{I})$, due to the non-linear transformation from the surrogate parameter space into the posterior space.
Instead, we aim to infer information about the location and structure of the posteriors $p(\theta\mid\tau^{(i)},D_{I})$ without resorting to expensive methods such as MCMC or requiring knowledge of the true parameter $\theta_{I}$. To this end, we draw samples $\{\theta_{I}^{(j)}\}_{j=1}^{J}\sim p(\theta_{I})$ from the prior. For each draw of surrogate parameters $\tau^{(i)}$, we compute the surrogate log-likelihood of the observed data $D_{I}$ as
@@ 3.1.1 Simulation Setup
$y=\beta_{0}+\beta_{1}x_{1}+\beta_{2}x_{2}+\dots+\beta_{p}x_{p}+\epsilon,\quad\epsilon\sim\mathcal{N}(0,1).$ (29)
@@
The model parameters $\mathbf{\beta}=(\beta_{0},\beta_{1},\dots,\beta_{p})$ were drawn from a $\mathcal{N}(0,1)$ distribution, while the covariates $x_{i}$ were generated with a mean of 0 and standard deviation sampled from a Gamma distribution with shape and rate parameters set to 10. The datasets were constructed using the R-package simstudy [49], enabling the generation of correlated data with a correlation coefficient of 0.3 between variables. For each dataset size, 20 datasets were created. Missing values were then introduced in $\lceil\pi\cdot N\rceil$ rows, with $\pi\in\{0.05,0.15,0.30\}$. Within each affected row, the target variable $y$ was always missing, along with $\lfloor\frac{p}{2}\rfloor$ randomly selected covariates.
@@
Each dataset was subsequently imputed $m=100$ times using the MICE algorithm, implemented in the R-package mice [20, 50]. The models where fitted on the imputed datasets via MCMC with the package R-package brms [51, 52, 53] with four chains for 2000 iterations each (including 1000 warm-up and 1000 sampling iterations). Two different prior specifications were tested: the default noninformative (flat) priors in brms and the regularized Horseshoe prior [54, 55], which is employed for regularization and model selection [56], running four chains for 2000 iterations each (including 1000 warm-up and 1000 sampling iterations).
@@
The simulation study evaluates four different methods. The first is a brute-force approach that fits each of the $m=100$ imputed datasets individually using MCMC. This computationally intensive method is implemented using brms. The second method employs the iterative algorithm described in Section 2.3, but omits the IWMM step. Specifically, after executing the PSIS step (Line 1 in Algorithm 1) and verifying its success for the given target distribution, the algorithm skips the IWMM step (Lines 1 to 1) and proceeds with a new proposal distribution. This variant is referred to as PSIS:single. The third method, termed PSIS:mixture, also relies solely on PSIS but utilizes a mixture distribution with $|I_{\rm mix}|=5$, as outlined in Algorithm 2 in Appendix B. The fourth method, introduced in Section 2.3, integrates both PSIS and IWMM – jointly referred to as PSIS+IWMM in the following. Preliminary simulations revealed that using a mixture proposal distribution within IWMM did not improve performance relative to a single proposal; in fact, it often degraded it. This reduction in performance is likely due to the close similarity in the structure of the target and single proposal distribution, whereas the use of a mixture distribution introduces greater divergence, thereby complicating the transformation step in IWMM. Therefore, no mixture distribution is tested for the PSIS+IWMM method.
@@
Performance Evaluation To evaluate the efficiency of the proposed method relative to the brute-force MCMC approach, we compare the number of log-likelihood evaluations. We do not assess the accuracy of the resulting marginalized posteriors, as we only accept importance resampling approximations from PSIS or IWMM when the corresponding shape parameters $\hat{k}$ or $\hat{k}_{\rm MM}$ fall below the threshold $k_{\rm threshold}$. This criterion already ensures that only reliable approximations of the target distribution are used [28], as exemplified in Figure 2.
The models for the PSIS and IWMM proposals are fitted in brms [51, 52] using Stan [53], where Hamiltonian Monte Carlo (HMC) [57], more specifically the no-U-turn-sampler (NUTS) [58], is employed, as the gold standard for accurate sampling. The majority of NUTS’s computational cost arises from the repeated evaluation of the log-probability and its gradient, referred to as gradient evaluations. The log-probability evaluations correspond to the total unnormalized log-posterior density, incorporating both the log-prior of the parameters and the log-likelihood of the observed data given the parameters. In contrast, gradient log-probability evaluations refer to computing the derivative of the log-probability function with respect to all model parameters. Evaluating these gradients is crucial for HMC, as it relies on using gradient information to navigate the parameter space effectively. Gradients are computed via automatic differentiation [59], and typically account for approximately 80% of the total computational time [60]. Therefore, the number of gradient evaluations is a key factor when comparing the computational costs of different methods. In addition, calculating the importance ratios for PSIS and IWMM requires log-likelihood evaluations – both for the numerator and denominator of the ratios. Due to the structure of the importance ratios, evaluating the log-prior terms separately is unnecessary. To compare computational effort across methods, we count log-likelihood evaluations from PSIS and IWMM together with the log-probability evaluations used in HMC. While this slightly overstates the cost for PSIS and IWMM – since log-likelihoods are usually cheaper to compute than full log-probabilities – it offers a consistent and conservative measure of computational demand.

### v2

@@ 2.3 Iterative Method
1:Variable of interest $\theta\sim p(\theta\mid C)$, underlying variable $\tau\sim p(\tau\mid C)$
2:Draw $m$ samples $\tau^{(1)},\dots,\tau^{(m)}\sim p(\tau\mid C)$
3:Set index set $I\leftarrow\{1,\dots,m\}$
4:while $|I|>0$ do
5:  Select representative value $\tau^{*}\in\{\tau^{(i)}\}_{i\in I}$ and run MCMC to access the posterior distribution through samples $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p(\theta\mid\tau^{*})$.
6:  Update index set II such that $\tau^{*}\notin\{\tau^{(i)}\}_{i\in I}$.
7:  for $i\in I$ do
8:   Run PSIS with target $p(\theta\mid\tau^{(i)})$, proposal $p(\theta\mid\tau^{*})$, obtain importance weights and metric $\hat{k}_{i}$
9:   if $\hat{k}_{i}<0.7$ then
10:     Use calculated importance weights and draws $\{\theta^{(s)}_{*}\}_{s=1}^{S}\sim p(\theta\mid\tau^{*})$ for importance    resampling and save the resampled draws as posterior draws corresponding to $p(\theta\mid\tau^{(i)})$.
11:     Remove $i$ from II: $I\leftarrow I\setminus\{i\}$
12:   else
13:     Run IWMM with target $p(\theta\mid\tau^{(i)})$ and proposal $p(\theta\mid\tau^{*})$     and gain importance weights and metric $\hat{k}_{\text{MM},i}$.
14:     if $\hat{k}_{\text{MM},i}<0.7$ then
15:      Use calculated importance weights and transformed draws $\{\breve{\theta}^{(s)}_{*}\}_{s=1}^{S}$ for importance      resampling and save the resampled draws as posterior draws corr. to $p(\theta\mid\tau^{(i)})$}.
16:      Remove $i$ from II: $I\leftarrow I\setminus\{i\}$
17:     end if
18:   end if
19:  end for
20:end while
21:Calculate the marginalized posterior via Monte Carlo Estimate $p(\theta\mid C)\approx\frac{1}{m}\sum_{i=1}^{m}p\left(\theta\mid\tau^{(i)}\right)$.
Algorithm 1 provides accurate posterior draws through MCMC or resampled posterior draws through PSIS or IWMM for all $\{\tau^{(i)}\}_{i=1}^{m}$, corresponding to the desired distribution $p(\theta\mid\tau^{(i)})$}. This allows the calculation of the Monte Carlo estimate from Equation (2), enabling us to approximate the marginalized posterior distribution $p(\theta\mid C)$. Since each posterior approximation used in the calculation is either based on computationally expensive MCMC or on PSIS/IWMM, where the quality of the approximation is ensured by accepting the resampling only when $\hat{k}$ or $\hat{k}_{\rm MM}$ is below the threshold, the accuracy of the estimate is maintained. The algorithm improves efficiency compared to running MCMC for all $m$ values separately, as each iteration (cycles of the while-loop in row 4) requires only one MCMC run for the proposal distribution, reducing the number of MCMC runs significantly. The improvement in efficiency is tested through extensive simulation studies in Section 3.
@@ 2.4 Special Case: Missing Data Problem
We now consider the setting in which the observed dataset $D_{\rm obs}=(y_{\rm obs},X_{\rm obs})$ contains missing values, potentially affecting both the target variable and the predictors. While there are straightforward approaches to handling missing data — such as listwise or pairwise deletion, and mean imputation [7] — these methods fail to propagate the uncertainty introduced by the missing values into the parameter estimates. A more principled approach that allows for uncertainty propagation is Multivariate Imputation using Chained Equations (MICE) [6]. MICE is an iterative algorithm that imputes the missing values in the dataset by generating $m$ completed versions of the data, each containing potentially different imputed values at the locations of the missing entries [78]. These multiple imputations help capture the uncertainty associated with the missing values in the original dataset $D_{\rm obs}$. We define the $m$ imputed datasets as $D^{(1)},\dots,D^{(m)}$, with observations in each dataset $D^{(i)}$ indexed by the superscript $(i)$ and denote the imputation process performed by MICE via the distribution $p_{\text{MICE}}(D\mid D_{\rm obs})$. Our goal is to characterize the marginalized posterior distribution of the model parameters $\theta$, which is obtained by integrating over the uncertainty in the imputed datasets:
$p_{\text{MICE}}(\theta\mid D_{\rm obs})=\int p$ $(\theta\mid D)\,p_{\text{MICE}}(D\mid D_{\rm obs})\,dD$ (27)
$\overset{\text{Monte Carlo}}{\approx}$ $\frac{1}{m}\sum_{i=1}^{m}p(\theta\mid D^{(i)}).$ (28)
For this problem setting, our method presented in Section 2.3 can be applied, with the context $C$ being the observed dataset $D_{\rm obs}$. A structural visualization of the iterative method (cf. Algorithm 1) is illustrated in Figure 3 for the special case of working with missing data..
@@
Calculation of Importance Ratios The specific case of handling missing data offers additional advantages that allow our algorithm to operate even more efficiently. In particular, the computation of importance ratios for PSIS and IWMM can be further optimized in this setting — an improvement not necessarily feasible in the more general iterative approach described in Section 2.3. This efficiency gain stems from the structural similarity among the imputed datasets $\{D^{(i)}\}_{i\in I}$, which differ only in the entries corresponding to originally missing values. We denote by $J^{*}\subset\{1,\dots,N\}$ the index set of observations where these differences occur.
@@
$r_{i}(\theta_{*}^{(s)})$ $=\frac{p(\theta_{*}^{(s)}\mid D^{(i)})}{p(\theta_{*}^{(s)}\mid D^{*})}=\frac{p(\theta_{*}^{(s)}\mid y^{(i)},X^{(i)})}{p(\theta_{*}^{(s)}\mid y^{*},X^{*})}$ (29)
$\propto\frac{p(y^{(i)}\mid\theta_{*}^{(s)},X^{(i)})p(\theta_{*}^{(s)})}{p(y^{*}\mid\theta_{*}^{(s)},X^{*})p(\theta_{*}^{(s)})}$ (30)
$=\frac{p(y^{(i)}\mid\theta_{*}^{(s)},X^{(i)})}{p(y^{*}\mid\theta_{*}^{(s)},X^{*})}\,,s=1,\dots,S.$ (31)
@@
$r_{i}(\theta_{*}^{(s)})$ $=\frac{\prod_{n=1}^{N}p(y_{n}^{(i)}\mid\theta_{*}^{(s)},X^{(i)})}{\prod_{n=1}^{N}p(y_{n}^{*}\mid\theta_{*}^{(s)},X^{*})}$ (32)
$=\frac{\prod_{n\in J^{*}}p(y_{n}^{(i)}\mid\theta_{*}^{(s)},X^{(i)})}{\prod_{n\in J^{*}}p(y_{n}^{*}\mid\theta_{*}^{(s)},X^{*})}\,$, (33)
$s=1,\dots,S$, because $p(y_{n}^{(i)}\mid\theta_{*}^{(s)},X^{(i)})={p(y_{n}^{*}\mid\theta_{*}^{(s)},X^{*})}$ for all $n\notin J^{*}$. This reduction allows us to evaluate only the likelihood terms that correspond to differing rows, decreasing the number of pointwise likelihood evaluations by a factor of $|J^{*}|/N$, where $|J^{*}|$ corresponds to the number of rows with missing values. This optimization significantly reduces the computational effort required to compute the importance ratios for PSIS and IWMM, thereby making the overall workflow more efficient.
@@ 2.5 Special Case: Surrogate Modeling
Inference Step Next, we consider a new set of observations — typically real-world measurements — denoted by $D_{I}=\{y_{I}^{(n)}\}_{n=1}^{N_{I}}$}. We assume that the (implicit) data-generating process underlying these observations is well captured by the simulator $\mathcal{M}$. In this context, the measurement data $D_{I}$ is thought to be generated by the simulator using some unknown input parameter $\theta_{I}$:
@@
$p(\theta_{I}\mid D_{T},D_{I})$ $=\int p(\theta_{I}\mid\tau,D_{I})\,p(\tau\mid D_{T})\,d\tau$ (38)
$\overset{\rm MC}{\approx}\frac{1}{m}\sum_{i=1}^{m}p(\theta_{I}\mid\tau^{(i)},D_{I})\,.$ (39)
@@
The selected draws should be representative in the sense that their corresponding posterior distributions collectively cover the diversity of possible posterior shapes — or, in the case of selecting a single draw, that the draw is central in terms of its posterior characteristics.
A straightforward approach is to apply clustering algorithms such as $k$-means [48] or $k$-medoids [38] in the space of the draws $\tau$. However, we cannot be sure that the similarity between surrogate parameter draws $\tau^{(i)}$ and $\tau^{(j)}$ does reliably reflect the similarity of their respective posterior distributions, $p(\theta_{I}\mid\tau^{(i)},D_{I})$ and $p(\theta_{I}\mid\tau^{(j)},D_{I})$, due to the non-linear transformation from the surrogate parameter space into the posterior space.
Instead, we aim to infer information about the location and structure of the posteriors $p(\theta_{I}\mid\tau^{(i)},D_{I})$} without resorting to expensive methods such as MCMC or requiring knowledge of the true parameter $\theta_{I}$. To this end, we draw samples $\{\theta_{I}^{(j)}\}_{j=1}^{J}\sim p(\theta_{I})$ from the prior. For each draw of surrogate parameters $\tau^{(i)}$, we compute the surrogate log-likelihood of the observed data $D_{I}$ as
@@
Figure 4 provides an exemplary visualization of the selection results. It shows the posterior distributions $p(\theta_{I}\mid\tau^{(i)},D_{I})$ for all surrogate realizations $\tau^{(i)}$ in the logistic-surrogate simulation study (Section 3.3). The $N_{\mathrm{mix}}=10$ distributions highlighted in blue are those selected by the proposed method. The selected posteriors are well distributed across the range of plausible values of $\theta_{I}$, indicating broader coverage than would typically be obtained from a random subset.
Figure 4: Posterior distributions $p(\theta_{I}\mid\tau^{(i)},D_{I})$ for $i=1,\dots,100$ from the logistic-surrogate simulation study in Section $3.$ The $N_{\mathrm{mix}}=10$ distributions highlighted in red were selected by the proposed method based on log-likelihoods evaluated over prior samples.
@@ 3.1.1 Simulation Setup
$y=\beta_{0}+\beta_{1}x_{1}+\beta_{2}x_{2}+\dots+\beta_{p}x_{p}+\epsilon,$ (42)
@@
$\epsilon\sim\mathcal{N}(0,1)$. The model parameters $\mathbf{\beta}={(\beta_{0},\beta_{1},\dots,\beta_{p})}$} were drawn from a $\mathcal{N}(0,1)$ distribution, while the covariates $x_{i}$ were generated with a mean of 0 and standard deviation sampled from a Gamma distribution with shape and rate parameters set to 10. The datasets were constructed using the R-package simstudy [27], enabling the generation of correlated data with a correlation coefficient of 0.3 between variables. For each dataset size, 20 datasets were created. Missing values were then introduced in $\lceil\pi\cdot N\rceil$ rows, with $\pi\in\{0.05,0.15,0.30\}$. Within each affected row, the target variable $y$ was always missing, along with $\lfloor\frac{p}{2}\rfloor$ randomly selected covariates.
Each dataset was subsequently imputed m=100{m=100} times using the MICE algorithm, implemented in the R-package mice [7, 8]. Choosing the number of imputations $m$ is nontrivial and depends on several factors such as the fraction of missing information, the desired stability of the resulting inference, and practical constraints in the application. Further guidance on selecting $m$ is provided in [28, 32]. In our setting, we examined how the posterior distributions depend on $m$ and on the number of samples drawn per imputed dataset $S$. Under the gold-standard approach — running MCMC separately for each imputed dataset — we observed no visible differences between the resulting posterior distributions for $m=10,S=4000$; $m=100,S=40$ and $m=1000,S=4$ in our case (see Figure 12 in Appendix C).
@@
The models were fitted on the imputed datasets via MCMC with the package R-package brms [9, 12, 76] with four chains for 2000 iterations each (including 1000 warm-up and 1000 sampling iterations). Two different prior specifications were tested: the default noninformative (flat) priors in brms and the regularized Horseshoe prior [16, 60], which is employed for regularization and model selection [15].
@@
The simulation study evaluates four different methods. The first is a brute-force approach that fits each of the $m=100$ imputed datasets individually using MCMC. This computationally intensive method is implemented using brms. The second method employs the iterative algorithm described in Section 2.3, but omits the IWMM step. Specifically, after executing the PSIS step (Line 8 in Algorithm 1) and verifying its success for the given target distribution, the algorithm skips the IWMM step (Lines 12 to 18) and proceeds with a new proposal distribution. This variant is referred to as PSIS:single. The third method, termed PSIS:mixture, also relies solely on PSIS but utilizes a mixture distribution with $|I_{\rm mix}|=5$, as outlined in Algorithm 2 in Appendix B. The fourth method, introduced in Section 2.3, integrates both PSIS and IWMM — jointly referred to as PSIS+IWMM in the following. Preliminary simulations revealed that using a mixture proposal distribution within IWMM did not improve performance relative to a single proposal; in fact, it often degraded it. This reduction in performance is likely due to the close similarity in the structure of the target and single proposal distribution, whereas the use of a mixture distribution introduces greater divergence, thereby complicating the transformation step in IWMM. Therefore, no mixture distribution is tested for the PSIS+IWMM method.
@@
Performance Evaluation To evaluate the efficiency of the proposed method relative to the brute-force MCMC approach, we compare the number of log-likelihood evaluations. The accuracy of the resulting marginalized posteriors should be satisfactory, as we only accept importance resampling approximations from PSIS or IWMM when the corresponding shape parameters $\hat{k}$ or $\hat{k}_{\rm MM}$ fall below the threshold $k_{\rm threshold}$. This criterion already ensures that only reliable approximations of the target distribution are used [77], as exemplified in Figure 2. To further validate this, we compute the mean absolute differences between posterior summaries obtained via MCMC and via the PSIS+IWMM approach, considering the posterior mean, standard deviation, and the 5% and 95% quantiles.
The models for the PSIS and IWMM proposals are fitted in brms [9, 12] using Stan [76], where Hamiltonian Monte Carlo (HMC) [53], more specifically the no-U-turn-sampler (NUTS) [33], is employed, as the gold standard for accurate sampling. The majority of NUTS’s computational cost arises from the repeated evaluation of the log-probability and its gradient, referred to as gradient evaluations. The log-probability evaluations correspond to the total unnormalized log-posterior density, incorporating both the log-prior of the parameters and the log-likelihood of the observed data given the parameters. In contrast, gradient log-probability evaluations refer to computing the derivative of the log-probability function with respect to all model parameters. Evaluating these gradients is crucial for HMC, as it relies on using gradient information to navigate the parameter space effectively. Gradients are computed via automatic differentiation [14], and typically account for approximately 80% of the total computational time [82]. Therefore, the number of gradient evaluations is a key factor when comparing the computational costs of different methods. In addition, calculating the importance ratios for PSIS and IWMM requires log-likelihood evaluations — both for the numerator and denominator of the ratios. Due to the structure of the importance ratios, evaluating the log-prior terms separately is unnecessary. To compare computational effort across methods, we count log-likelihood evaluations from PSIS and IWMM together with the log-probability evaluations used in HMC. While this slightly overstates the cost for PSIS and IWMM — since log-likelihoods are usually cheaper to compute than full log-probabilities — it offers a consistent and conservative measure of computational demand.